Simplify. (t + s^2 )(t^2 – s) Simplify. (t + s^2 )(t^2 – s)

Simplify. 
 (t + s^2 )(t^2 – s)
Simplify. 
 (t + s^2 )(t^2 – s)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(t + s^2)(t2–s)
= t (t^2–s)+s^2(t^2–s)
= t^3 –st + s^2t^2 –s^3
= t^3–s^3–st + s^2t^2

Thanks